OpenStudy (anonymous):

What is the direct variation equation for the above table, using t for sales tax and p for price? t = p + 0.06 t = 0.06p t = t =

OpenStudy (jdoe0001):

the table above? you must have posted above the screen, since it's not very noticeable

OpenStudy (jdoe0001):

posted it rather... anyhow.. got table?

OpenStudy (anonymous):

t = p + 0.06 t = 0.06p t = t =

OpenStudy (jdoe0001):

well, lets tak a look at ... hmm say the 3rd value for example \(\begin{array}{ccllll} price&sales\ tax\\ (x)&(y) \\\hline\\ 1.00&0.06\\ 2.00&0.12\\ 3.00&0.18&\impliedby this\ one\\ 4.00&0.24 \end{array} \\ \quad \\ \begin{array}{cccccclllll} \textit{something}&&\textit{varies directly to}&&\textit{something else}\\ \quad \\ \textit{something}&=&{\color{brown}{ \textit{some value}}}&\cdot &\textit{something else}\\ \quad \\ y&=&{\color{brown}{ n}}&\cdot&x \\\hline\\ && y={\color{brown}{ n }}x \end{array}\\ \quad \\ \\ \quad \\ y = 0.18\qquad x=3.00\qquad thus\implies 0.18={\color{brown}{ n}}\cdot 3.00 \\ \quad \\ \textit{solve for }{\color{brown}{ n}}\ or\ {\color{brown}{ constant\ of\ variation}}\)

OpenStudy (jdoe0001):

what would that give you for "n"?

OpenStudy (anonymous):

i think its 0.54 if i did it the right way

OpenStudy (jdoe0001):

well... how did you get 0.54 anyway?

OpenStudy (anonymous):

0.18 x 3.00

OpenStudy (jdoe0001):

hmmm to get "n", you'd want to divide both sides by 3.00

OpenStudy (anonymous):

0.06

OpenStudy (jdoe0001):

yeap \(\bf y = 0.18\qquad x=3.00\qquad thus\implies 0.18={\color{brown}{ n}}\cdot 3.00 \\ \quad \\ \cfrac{0.18}{3.00}={\color{brown}{ n}}\implies 0.06=n\qquad thus \\ \quad \\ y={\color{brown}{ n}}x\implies y={\color{brown}{ 0.06}}x\)

OpenStudy (jdoe0001):

well, in this case.. price (p) is the "x" and the tax(t) is the 'y" or dependent so, that'll make it \(\bf t={\color{brown}{ 0.06}}p\)

OpenStudy (anonymous):

Find the constant of variation, k, for the above table. What does this represent? how would i find this out using the same table we did? i just need steps

OpenStudy (jdoe0001):

heheh well.. read above what "n: is :)

OpenStudy (jdoe0001):

\(y = 0.18\qquad x=3.00\qquad thus\implies 0.18={\color{brown}{ n}}\cdot 3.00 \\ \quad \\ \textit{solve for }{\color{brown}{ n}}\ or\ {\color{brown}{ constant\ of\ variation}}\)

OpenStudy (jdoe0001):

if you want to make it a "k" then, so \(y = 0.18\qquad x=3.00\qquad thus\implies 0.18={\color{brown}{ k}}\cdot 3.00 \\ \quad \\ \textit{solve for }{\color{brown}{ k}}\ or\ {\color{brown}{ constant\ of\ variation}}\)

OpenStudy (anonymous):

oh so theres no differences its just the variables that are

OpenStudy (jdoe0001):

well, what you found,when solving for "n", was the "constant of variation", that's what "n" is, and n = 0.06 :) or k if you wish to call it that

OpenStudy (anonymous):

oh okay thank you

OpenStudy (jdoe0001):

yw